The following activities are part of a project to be scheduled using CPM:

ACTIVITY IMMEDIATE PREDECESSOR TIME (WEEKS)
A 3
B A 9
C A 5
D C 2
E B D 7
F D 6
G E F 4

b. What is the critical path?

A-B-D-F-G
A-B-E-G
A-C-D-F-G
A-C-D-E-G

c. How many weeks will it take to complete the project?

Number of weeks            

d. How much slack does activity B have?

Slack of activity B             week(s)
Answer #1
step 1 find early start, and early finish from the start of the project for each activity
step 2 Find Maximum EF and keep it as the LF of last activities. Max of EF's is 23
step 3 calculate LS for each activity after calculating its EF
Activity Duration, D Early start, ES=Max of early finish of preceding activities Early finish, EF = ES + D Late finish, LF= Min of LS of successor activities Late start, LS= LF - D Total slack= LF-EF Critical activity, activities with 0 slack time
A 3 0 3 3 0 0 Yes
B 9 3 12 12 3 0 Yes
C 5 3 8 10 5 2 No
D 2 8 10 12 10 2 No
E 7 12 19 19 12 0 Yes
F 6 10 16 19 13 3 No
G 4 19 23 23 19 0 Yes

b. Hence the critical path is A-B-E-G

c. it will take 23 weeks

d. slack of B is 0.